What is the formula to calculate the length of the diagnal side of a right angle triangle if the other 2 dimensions are known?
Answer by thechamp1011(19)

Use the pythegorean theorem to solve the length of the diagonal side of a right triangle if the other two side of the triagle is know.
EX# Draw a right triangle and assume the diagonal is AC (unknown)
and the two sides known are AB=2 and BC=3
Using the pythegorean theorem: AB^2+BC^2=AC^2
2^2+3^2=AC^2
4+9=AC^2
13=AC^2
Therefore, AC=3.61.....by taking the sqrt(13)